# Word Alignment as Preference for Machine Translation

### v1

@@ 4.3 The design of evaluation
We focus on optimizing LLM-based MT models to avoid hallucination and omission. However, to our best knowledge, there is no benchmark measuring MT models specifically for this issue, making the evaluation very challenging. Improving the BLEU or COMET score does not necessarily mean reducing hallucination and omission because there are other factors such as mistranslation and fluency. In addition, hallucination is relatively infrequent, although very severe once happens. To intuitively validate whether our approach is capable of mitigating hallucination and omission in MT, we design several evaluation strategies in this section.
@@
We first select instances that the baseline model does not perform well on. This subset of instances is labeled as hard instances in this paper. The subset of the remaining examples is labeled as easy instances. Specifically, $N$ instances with the lowest COMET score are selected from the test set for each translation direction. As hard examples tend to include more hallucination and omission, we report the comparison of models on hard examples and remaining examples, respectively. In the experiment, we sample three subsets where $N=100$, $N=200$ and $N=500$. The experimental analysis can be found in §5.1. Note that the hard instances are only selected for evaluation. We do not differentiate hard or easy instances in the training set. Only word alignment signal is used to select preferred dataset for a fair comparison.
@@
Table 1: Average coverage score calculated by GPT-4 for different level of hallucination or omission. The Pearson Correlation between the annotated labels and GPT-4 coverage scores is also reported.
@@
Besides the BLEU and COMET in hard instances, a direct estimate of the degree of hallucination and omission in translation is still needed. As we mentioned earlier that improving the BLEU and COMET score does not necessarily mean reducing hallucination and omission because there are other factors such as mistranslation and fluency, we utilize the generalization and reasoning ability of LLM (Kojima et al., 2022; Mitchell et al., 2023; Wei et al., 2023) to achieve this direct evaluation. We use one of the most powerful LLM, GPT-4121212https://openai.com/research/gpt-4, as the evaluator. LLM is prompted to check whether the given translation has hallucination or omission referring to the given source texts. A coverage score between 0 and 100 is output as the degree metric. The prompt used is shown in Figure 5.
@@
Table 1 shows the average score of the degree of coverage predicted by GPT-4. The examples from HalOmi are split into three subsets based on the labels. We merged the “Partial hallucination and omission” and “Small hallucination and omission” in the original because the number of examples in these two categories is small. It clearly demonstrates that examples annotated as “No hallucination and omission” have a higher coverage score predicted by GPT-4 and those in “Full hallucination and omission” have an extremely low coverage score. As a result, using GPT-4 is an effective way to assess whether a translation has the problem of hallucination or omission.
@@
Table 3: Human evaluation on “zh-en” when N=100. Translation quality is the ratio of examples where the corresponding model generates better translation in general. The remaining columns present the ratio of examples in which the corresponding degree of hallucination or omission occurs. Better performance is highlighted with bold fonts.
@@ 5.1 Evaluation on hard instances
In §4.3 we introduce how to select hard instances from the test set and explain why hard instances are suitable to assess hallucination and omission. In this section, we evaluate our model on these hard instances and the remaining examples, respectively. Figure 6 demonstrates the results when the number of sampled instances $N=100,200$, and $500$, respectively. The following findings can be concluded:
  • Our model consistently outperforms the baseline on hard instances in most translation directions, for both BLEU and COMET metrics.
  • Our model reaches competitive performance compared to the baseline for both BLEU and COMET.
  • With increasing the number of sampled hard instances, the improvement gained by our model gets smaller.
These results indicate that WAP mitigates hallucination and omission to a certain extent, because these issues are more likely to occur in hard instances. In addition, with the improvement in the hard instances, our model remains competitive to the baseline in the remaining easy instances. It is reasonable that there is no significant difference in the easy instances because the compared models are generally good. The challenging part should be in the hard ones. Moreover, it is also observed that with increasing $N$, the improvement gets narrower. The reason is that more relatively easy instances are included in the subset. This is another evidence that WAP provides gains particularly for hallucination and omission in MT. The specific numeric results and the overall results for the entire test set are shown in §B.
@@ 5.2 Direct evaluation of hallucination and omission by GPT-4
In addition to improving hard examples, which is more likely to have hallucination and omission, direct evaluations of them are also needed to confirm the effectiveness of the proposed WAP. In §4.3 we have verified the usefulness of GPT-4 as an evaluator with experiments. In this section, we prompt GPT-4 to directly predict a coverage score as the metric of hallucination and omission. The results are demonstrated in Table 2. The reported number is the average of the coverage scores in hard examples. The results show that our model outperforms the baseline in all translation directions except $en\leftrightarrow is$. Specifically in the average score of all translation directions, WAP outperforms the baseline model by 4.96, 1.63 and 1.24 when N=100, 200 and 500, respectively. The trend is similar to that of §5.1, which directly indicates that the LLM-based MT model is steered to avoid generating hallucination and omission in MT with the preference dataset we constructed.
@@ 5.3 Human evaluation
Although the validity of GPT-4 as evaluator for hallucination and omission has been demonstrated in §4.3 and Table 1, we conduct a human evaluation to further verify our findings, as LLM could still be unreliable. The subset of “N=100” on “zh-en” is selected. Three volunteers who speak Chinese and English are asked to assess the quality of the translation and the degree of hallucination and omission for the baseline and our model, without knowing which model generates the translations. Table 3 demonstrates the results. In general, our model generates better translation in 39.66% of the examples, while the percentage for ALMA is 11.33%. Furthermore, it is observed that with DPO on word-alignment preferred data fine-tuning, the degree of both hallucination and omission decreases. Specifically, the percentage of “no hallucination” increases from 64% to 75.66%, and that of “small, partial, and full hallucination” decreases accordingly. The decrease in omission is more distinct, in which the percentage of “no omission” increase by 24%. Notably, for both hallucination and omission, the percentage of “full hallucination and omission” has decreased to 0 for our model. These results indicate that omission is more frequent than hallucination, and WAP can mitigate hallucination and omission in LLM-based MT model like ALMA to some extent.
@@ 5.4 Ablation study
Figure 7: Ablation study. Results in BLEU is demonstrated. Higher BLEU is better. For fair comparison the range of y-axis are the same for hard instances and easy instances. The result in COMET is in the same trend, which can be found in Figure 8 in the Appendix.
@@
Does the preferred data really better contribute to the training?
It is observed that FT_prefer significantly outperforms FT_reject in both hard and easy instances. This indicates that our proposed pipeline ensures that the samples are selected, leading to better translation quality.
Is the DPO preference tuning necessary?
Particularly, the filled area demonstrates the necessity of preference tuning using DPO. In hard instances FT_prefer can reach a competitive performance with a small gap. However, in easy instances, FT_prefer largely underperforms WAP and ALMA, which limits the practicality of it. The possible reason for the different performance in the hard and easy instances is the direct fine-tuning. Directly fine-tuning on the preferred data without the comparison with rejected examples could cause a hard fitting to the word-aligned preference but ignore the general translation quality.
@@ 6 Conclusion
The problem of hallucination and omission, a long-standing problem in MT, could become more severe when an LLM is used because an LLM itself could hallucinate or omit in nature. In this paper, our aim is to mitigate this problem in LLM-based MT by optimizing the model toward a preference for better word alignment. We construct preference datasets by collecting translations using multiple MT tools and selecting the preferred translation with a higher coverage score output by a word aligner. DPO is then utilized to optimize the model towards the word-aligned preference. As evaluation of hallucination and omission is challenging, we design experiments that include selecting hard instances and using GPT-4 to directly predict coverage score, ensuring an effective evaluation. The experiments demonstrate that the proposed WAP mitigates hallucination and omission in ten translation directions, especially in hard instances.
@@
The primary limitation of our method stems from the imperfections of the word alignment model. Within our approach, it is inevitable to encounter some alignment errors, which we address through a filtering method. However, this solution adds complexity and clutter to the method. Additionally, the effectiveness of our method is diminished for low-resource language translations due to the limited number of parallel sentences available. Lastly, our reliance on the GPT-4 API to evaluate the results introduces a significant cost factor. We aim to find a cost-free alternative for this evaluation process in future work.
@@
All datasets and checkpoints used in this paper are copyright-free for research purposes. Previous studies are properly cited and discussed. This research aims to improve LLM-based machine translation models with word alignment preference data, and the preference is made by an automatic word aligner. We do not introduce additional bias to particular communities.
Appendix A Example analysis
A.1 Examples of the preference dataset
@@
source {CJK*} UTF8gbsn“我想,在考虑重播时,可以解决这个问题”,Coker 说道。
chosen (gpt-3.5) "I think, when considering replay, this issue can be resolved," Coker said.
rejected (human) "$<$I think that when I think about$>$ the replay, $<$I think that$>$ we can probably work it out," Coker said.
Example 2 (Chinese-English)
source {CJK*} UTF8gbsn<<<富勒>>>在政变图谋失败后
chosen (deepl) $<$Fuller$>$ after the failed coup attempt
rejected (human) After the failure of the attempted coup,
Example 3 (English-Chinese)
chosen (gpt-3.5) {CJK*} UTF8gbsn <<<最初是一个一居室的房产,布局错综复杂 - 你必须穿过厨房才能到达卧室>>> - 然而乔安妮想要增加存储空间和一个夹层,以充分利用宽敞的天花板高度。
rejected (deepl) {CJK*} UTF8gbsn 乔安妮希望增加储藏空间和一个夹层,充分利用宽敞的天花板高度。
Table 4: Examples in the preference dataset. The hallucination in rejected examples and omission in the source sentence are highlighted with <<<>>>. The corresponding contents that are omitted in the rejected example are highlighted with <<<>>> in the chosen example.
Table 4 includes three examples in our dataset, in which the source sentence, the chosen and rejected translations are shown. Refer to §3.4 for a detailed construction of the dataset. Example 1: the rejected translation is from human annotation, in which it repeats the term of “I think” unnaturally. The possible reason could be the resource of the parallel data, e.g., direct collection from transcriptions. Example 2: “Fuller” is omitted by human annotation while translated by DeepL. Example 3: the chosen translation is from gpt-3.5-turbo that completely translates the source sentence. In contrast, the translation by DeepL omits the first half.
A.2 Translation examples
Example 1 (English-Chinese)
Source Sunday Best: Enter 1880s New York $<$in HBO’s "The Gilded Age"$>$
Translation (Baseline) {CJK*} UTF8gbsn周日最佳:进入 1880 年代的纽约
Translation (Ours) {CJK*} UTF8gbsn周日最佳:进入 1880 年代的纽约 <<<,在 HBO 的《金碧辉煌时代》>>>
Example 2 (English-Chinese)
Translation (Baseline) {CJK*} UTF8gbsn粘贴和悬挂<<<卡扣的内部卡扣用于保持卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣>>>
Translation (Ours) {CJK*} UTF8gbsn内固定和悬挂标签内固定和悬挂标签用于保持薄膜在位,与我们提供的标签对应。
Example 3 (Chinese-English)
Source {CJK*} UTF8gbsn不知道要<<<等到什么时候>>>
Translation (Baseline) I don’t know when
Translation (Ours) I don’t know $<$how long I have to wait$>$
Table 5: Translation Examples. The hallucination in translation by the baseline and the omission in the source sentence are highlighted with <<<>>>. The corresponding contents that are omitted from the baseline are highlighted with $<$ $>$ in our translation.
@@ A.2 Translation examples
(a) Hard instances
(b) Easy instances
Figure 8: Ablation study. Results in COMET is demonstrated. Higher COMET is better. For fair comparison the range of y-axis are the same for hard instances and easy instances. Refer to §5.4 for discussion.
Appendix B Specific results

### v2

@@ 5.2 The design of evaluation
We focus on optimizing LLM-based MT models to avoid hallucination and omission. However, to our best knowledge, there is no benchmark measuring MT models specifically for this issue, making the evaluation very challenging. Improving the BLEU or COMET score does not necessarily mean reducing hallucination and omission because there are other factors such as mistranslation and fluency. In addition, hallucination is relatively infrequent, although very severe once it does occur. To intuitively validate whether our approach is capable of mitigating hallucination and omission in MT, we design several evaluation strategies in this section.
@@
We first select instances that the baseline model does not perform well on. This subset of instances is labeled as hard instances in this work. The subset of the remaining examples is labeled as easy instances. Specifically, $N$ instances with the lowest COMET score are selected from the test set for each translation direction. As hard examples tend to include more hallucination and omission, we report the comparison of models on hard examples and remaining examples, respectively. In the experiment, we sample three subsets where $N=100$, $N=200$ and $N=500$. The experimental analysis can be found in §6.1. Note that the hard instances are only selected for evaluation. We do not differentiate hard or easy instances in the training set. Only word alignment signal is used to select preferred dataset for a fair comparison.
Figure 6: Prompt to calculate the coverage score.
@@
Table 1: Average coverage score calculated by GPT-4 for different level of hallucination or omission. The Pearson Correlation between the annotated labels and GPT-4 coverage scores is also reported. Ideally, higher score should correlate to less hallucination and omission.
@@
Besides the BLEU and COMET in hard instances, a direct estimate of the degree of hallucination and omission in translation is still needed. As we mentioned earlier that improving the BLEU and COMET score does not necessarily mean reducing hallucination and omission because there are other factors such as mistranslation and fluency, we utilize the generalization and reasoning ability of LLM (Kojima et al., 2022; Mitchell et al., 2023; Wei et al., 2023) to achieve this direct evaluation. We use one of the most powerful LLM, gpt-4-0613101010https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4, as an evaluator. LLM is prompted to check whether the given translation has hallucination or omission referring to the given source texts. A coverage score between 0 and 100 is output as the degree metric. The prompt used is shown in Figure 6.
@@
Table 1 shows the average score of the degree of coverage predicted by GPT-4. The examples from HalOmi are divided into three subsets according to the labels. We merged the “Partial hallucination and omission” and “Small hallucination and omission” in the original because the number of examples in these two categories is small. It clearly demonstrates that examples annotated as “No hallucination and omission” have a higher coverage score predicted by GPT-4 and those in “Full hallucination and omission” have an extremely low coverage score. As a result, using GPT-4 is an effective way to assess whether a translation has the problem of hallucination or omission.
@@
Table 3: Human evaluation on “zh-en” when N=100. Translation quality is the measured by ratio of examples where WAP beats the baseline. The remaining columns present the ratio of examples in which the corresponding degree of hallucination or omission occurs. Better model is highlighted with bold fonts.
@@ 6.1 Evaluation on hard instances
In §5.2 we introduce how to select hard instances from the test set and explain why hard instances are suitable to assess hallucination and omission. In this section, we evaluate our model on these hard instances and the remaining examples, respectively. Figure 5 demonstrates the results when the number of hard instances $N=100,200$, and $500$, respectively. The following findings can be concluded:
  • WAP consistently outperforms the baseline in hard instances in most translation directions, for both BLEU and COMET metrics.
  • WAP generally reaches comparable performance compared to baseline for both BLEU and COMET.
  • With increasing the number of hard instances, the improvement gained by WAP decreases.
These results indicate that WAP mitigates hallucination and omission to a certain extent, because these issues are more likely to occur in hard instances. In addition, our model also remains competitive to the baseline in the remaining easy instances. It is reasonable that there is no significant difference because the compared models are generally good. The challenging part should be in the hard ones. Moreover, it is observed that with increasing $N$, the improvement gets narrower. The reason is that more relatively easy instances are included in the subset. This is another evidence that WAP provides gains particularly for hallucination and omission in MT. The specific numeric results and the overall results for the entire test set are shown in §C.
@@ 6.2 Direct evaluation of hallucination and omission by GPT-4
In addition to improving BLEU and COMET in hard examples prone to hallucination and omission, direct evaluations are also necessary to confirm the effectiveness of WAP. In §5.2 we have verified the usefulness of GPT-4 as an evaluator with experiments. In this section, we prompt GPT-4 to directly predict a coverage score as a metric for hallucination and omission. The results are demonstrated in Table 2. The reported number is the average of the coverage scores in hard examples. The results show that WAP outperforms the baseline in all directions except $en\leftrightarrow is$. In the overall average score across all translation directions, WAP outperforms the baseline model by 4.96, 1.63 and 1.24 when N=100, 200 and 500, respectively. The trend is similar to that in §6.1, directly indicating that the LLM-based MT model avoids hallucination and omission with the word-aligned preference.
@@ 6.3 Human evaluation
Although the validity of GPT-4 as evaluator for hallucination and omission has been demonstrated in §5.2 and Table 1, we conduct a human evaluation to further verify our findings, as LLM could still be unreliable. The subset of “N=100” on “zh-en” is selected. Three volunteers who speak Chinese and English are asked to assess the quality of the translation and the degree of hallucination and omission for the baseline and our model, without knowing which model generates the translations. Table 3 demonstrates the results. In general, our model generates better translation in 39.66% of the examples, while the percentage for ALMA is 11.33%. Furthermore, it is observed that with DPO on word-alignment preferred data fine-tuning, the degree of both hallucination and omission decreases. Specifically, the percentage of “no hallucination” increases from 64% to 75.66%, and that of “small, partial, and full hallucination” decreases accordingly. The decrease in omission is more distinct, in which the percentage of “no omission” increase by 24%. Notably, for both hallucination and omission, the percentage of “full hallucination and omission” has decreased to 0 for our model. These results indicate that omission is more frequent than hallucination, and WAP can mitigate them in the LLM-based MT model.
@@ 6.4 Ablation study
(c) Hard instances
(d) Easy instances
Figure 7: Ablation study. Results in BLEU is demonstrated. Higher BLEU is better. For fair comparison the range of y-axis are the same for hard instances and easy instances.
@@
Does the preferred data truly contribute more to training?
It is observed that FT_prefer significantly outperforms FT_reject in both hard and easy instances. This suggests that WAP effectively selects samples, improving translation quality. It highlights the importance of selecting high-quality training data, as even human-annotated data can be flawed Xu et al. (2024a).
Is DPO preference tuning necessary?
The filled area highlights the necessity of preference tuning with DPO. While FT_prefer performs competitively in hard instances, it significantly underperforms WAP and ALMA in easy instances, limiting its practicality. The possible reason for the different performance in the hard and easy instances can be the direct fine-tuning, which focusing solely on preferred data without comparing it to rejected examples can lead to overfitting to word-aligned preferences, neglecting overall translation quality.
@@ 7 Conclusion
The problem of hallucination and omission, a long-standing problem in MT, could become more severe when an LLM is used because an LLM itself could hallucinate or omit in nature. In this paper, our aim is to mitigate this problem in LLM-based MT by optimizing the model toward a preference for better word alignment. We construct preference datasets by collecting translations using multiple MT tools and selecting the preference pair with a higher coverage score output by a word aligner. DPO is then utilized to optimize the model towards the word-aligned preference. As evaluation of hallucination and omission is challenging, we design experiments that include selecting hard instances and using GPT-4 to directly predict coverage score, ensuring an effective evaluation, which indicates that the proposed WAP mitigates hallucination and omission, especially in hard instances. On the other hand, although WAP shows promise in mitigating hallucination and omission, the overall performance of MT in different language directions remains mixed, with slight increases in BLEU and decreases in COMET.
@@
The first limitation of our method stems from the imperfections of the word alignment model. Within our approach, it is inevitable to encounter some alignment errors, which we address through a filtering method. However, this solution adds complexity and clutter to the method. Additionally, the effectiveness of our method is diminished for low-resource language translations due to the limited number of parallel sentences available. From the perspective of experiments, we only evaluate the methods in English-centric translation pairs due to the lack of Non-English data, in which hallucination and omission could happen more frequently. In particular, the WMT-2024 General Machine Translation Task Kocmi et al. (2024) has adopted non-English language pairs, such as Czech-to-Ukrainian and Japanese-to-Chinese, which could expand our work in the future. Moreover, our reliance on the GPT-4 API to evaluate the results introduces a significant cost factor. In future work, our objective is to find a cost-free alternative to this evaluation process. Lastly, although WAP shows promise in mitigating hallucination and omission, the overall performance of MT in different language directions remains mixed, with slight increases in BLEU and decreases in COMET.
@@
All datasets and checkpoints used in this paper are copyright-free for research purposes. Previous studies are properly cited and discussed. This research aims to improve LLM-based machine translation models with word alignment preference data, and the preference is made by an automatic word aligner. We do not introduce additional bias to particular communities. We have obtained the consent of the annotation volunteers for this study.
Appendix A Experimental setup
Example 1 (Chinese-English) Coverage Score
source {CJK*} UTF8gbsn“我想,在考虑重播时,可以解决这个问题”,Coker 说道。
chosen (gpt-3.5) "I think, when considering replay, this issue can be resolved," Coker said. 94.03
rejected (human) "$<$I think that when I think about$>$ the replay, $<$I think that$>$ we can probably work it out," Coker said. 79.87
Example 2 (Chinese-English) Coverage Score
source {CJK*} UTF8gbsn<<<富勒>>>在政变图谋失败后
chosen (deepl) $<$Fuller$>$ after the failed coup attempt 83.76
rejected (human) After the failure of the attempted coup, 59.59
Example 3 (English-Chinese) Coverage Score
chosen (gpt-3.5) {CJK*} UTF8gbsn <<<最初是一个一居室的房产,布局错综复杂 - 你必须穿过厨房才能到达卧室>>> - 然而乔安妮想要增加存储空间和一个夹层,以充分利用宽敞的天花板高度。 83.76
rejected (deepl) {CJK*} UTF8gbsn 乔安妮希望增加储藏空间和一个夹层,充分利用宽敞的天花板高度。 69.97
Table 4: Examples in the preference dataset. The hallucination in rejected examples and omission in the source sentence are highlighted with <<<>>>. The corresponding contents that are omitted in the rejected example are highlighted with <<<>>> in the chosen example. The coverage is calculated by word aligner, refer to §3 for details.
Appendix B Example analysis
B.1 Examples of the preference dataset
Table 4 includes three examples in our dataset, in which the source sentence, the chosen and rejected translations are shown. Refer to §4 for a detailed construction of the dataset. Example 1: the rejected translation is from human annotation, in which it repeats the term of “I think” unnaturally. The possible reason could be the resource of the parallel data, e.g., direct collection from transcriptions. Example 2: “Fuller” is omitted by human annotation while translated by DeepL. Example 3: the chosen translation is from gpt-3.5-turbo that completely translates the source sentence. In contrast, the translation by DeepL omits the first half.
Example 1 (English-Chinese) Coverage Score
Source Sunday Best: Enter 1880s New York $<$in HBO’s "The Gilded Age"$>$
Translation (Baseline) {CJK*} UTF8gbsn周日最佳:进入 1880 年代的纽约 70.0
Translation (Ours) {CJK*} UTF8gbsn周日最佳:进入 1880 年代的纽约 <<<,在 HBO 的《金碧辉煌时代》>>> 100.0
Example 2 (English-Chinese) Coverage Score
Translation (Baseline) {CJK*} UTF8gbsn粘贴和悬挂<<<卡扣的内部卡扣用于保持卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣卡扣>>> 0.0
Translation (Ours) {CJK*} UTF8gbsn内固定和悬挂标签内固定和悬挂标签用于保持薄膜在位,与我们提供的标签对应。 60.0
Example 3 (Chinese-English) Coverage Score
Source {CJK*} UTF8gbsn不知道要<<<等到什么时候>>>
Translation (Baseline) I don’t know when 90.0
Translation (Ours) I don’t know $<$how long I have to wait$>$ 100.0
Example 4 (English-Chinese)
Source $<$Pictures come in$>$ a sturdy carded envelope $<$in the box with the frame.$>$
Translation (Baseline) {CJK*} UTF8gbsn画框附带一个坚固的卡片盒。
Translation (Ours) {CJK*} UTF8gbsn画框<<<在盒子里,图片>>>放在一个坚固的卡片盒里。
@@ B.1 Examples of the preference dataset
Source {CJK*} UTF8gbsn这个<<<锅>>>买得好值。
Translation (Baseline) It’s a good buy.
Translation (Ours) This $<$pot$>$ is worth buying.
Table 5: Translation Examples. The hallucination in translation by the baseline and the omission in the source sentence are highlighted with <<<>>>. The corresponding contents that are omitted from the baseline are highlighted with $<$ $>$ in our translation. The coverage is calculated by GPT-4, refer to §5.2 for details.
B.2 Translation examples
@@
On the other hand, WAP could also fail in some cases. Example 4: Although the baseline omits “pictures” and “box,” which our model successfully translates, the translation of our model is not fully correct. The source is “in the box with frame,” but our model’s translation is “{CJK*}UTF8gbsn画框在盒子里 (frame in the box).” Example 5: Although our model translates “pot ({CJK*}UTF8gbsn锅)” that is omitted by the baseline, the meaning of the sentence is incorrect. The source means "This pot is a good buy," but our translation is "This pot is worth buying." In general, our model performs well in terms of coverage, which is more related to hallucination and omission; however, the translation quality does not necessarily improve accordingly. The study of preference signals for both overall translation quality and reducing hallucination and omission is worth exploring.
Appendix C Overall MT Performance